An airplane travels 500m down the runway before taking off it starts from rest , moves with constant acceleration and becomes airborne after 20.0s. What is the speed when it takes off?
1
Expert's answer
2018-09-13T12:12:08-0400
s=1/2 at^2 ⇨ =a=2s/t^2 .
Substitute a for 2s/t^2 below:
v=√2as=√(2·2s/t^2 ·s)=2s/t=(2·500)/20=50 m/s.